# The Complexity of Resilience Problems via Valued Constraint Satisfaction

### v5

@@ 8.3. Regular Path Queries
We introduce two-way regular path queries and show that they have finite duals. A two-way regular path query (2RPQ) over a finite relational signature $\tau$ is a regular expression over the alphabet $\Sigma_{\tau}$ that consists of the symbols $R$ and $R^{-}$ for every binary relation symbol $R\in\tau$. We assume that regular expressions are built up from alphabet symbols, $\emptyset$, $\varepsilon$, $+$ (union), ; (concatenation), and $\cdot^{*}$ (Kleene star). As usual, we may omit the concatenation symbol for brevity.
@@
We consider a Boolean version of the resilience problem for 2RPQs. For a 2RPQ $\mu$ over a signature $\tau$ and a set $\sigma\subseteq\tau$ of exogeneous relations, this problem is as follows: given a finite (bag) database ${\mathfrak{A}}$ and a $k\geq 1$, decide whether it suffices to remove at most $k$ tuples from ${\mathfrak{A}}$ to achieve that $\mu({\mathfrak{A}})=\emptyset$. It is folklore that given a 2RPQ $\mu$ and a database ${\mathfrak{A}}$, one can decide in polynomial time whether $\mu({\mathfrak{A}})=\emptyset$, e.g. by forming the product of ${\mathfrak{A}}$ and a finite automaton $A$ with $L(A)=L(\mu)$ and then checking whether some element with a final state in the second component is reachable from any element with the initial state in the second component. Consequently, for any 2RPQ $\mu$ with exogeneous relations from any set $\sigma$ the resilience problems is in NP by a straightforward guess-and-check algorithm.
@@
Let $\mu$ be a 2RPQ over a relational signature $\tau$ and let $\sigma\subseteq\tau$. Then the resilience problem for $\mu$ with exogenous relations from $\sigma$ is in P or NP-complete. Moreover, it is decidable whether the resilience problem for a given 2RPQ $\mu$ and set of exogeneous relations $\sigma$ is in P.
@@
Consider the 2RPQ $\mu$ over the signature $\{R\}$ given by $R(RR^{-})^{*}RR$. It is well-known (see (Hell and Nešetřil, 2004)) that the relational $\{R\}$-structure ${\mathfrak{B}}$ with domain $\{0,1,2\}$ and the relation $R^{{\mathfrak{B}}}:=\{(0,1),(1,2)\}$ is a dual of $\mu$. Let $\Gamma:=\Gamma({\mathfrak{B}};\emptyset)$. Note that the crisp unary relation $\{0,1\}$ is expressible in $\Gamma$ by
@@
Let $R^{\prime}(x,y):=R(x,y)+\phi(x)+\phi(y)$ and note that $R^{\prime}\in\langle\Gamma\rangle$. It is easy to see that the valued structure ($(\{0,1,2\};R^{\prime})$ is homomorphically equivalent to the valued structure $\Gamma_{<}$ from Example 2.4. In other words, $\Gamma$ pp-constructs $\Gamma_{<}$. Since $\operatorname{VCSP}(\Gamma_{<})$ is NP-complete (Example 4.7), $\operatorname{VCSP}(\Gamma)$ is NP-complete by Lemma 5.13 (every finite-domain VCSP is in NP). Therefore, the resilience problem for the 2RPQ $\mu$ is NP-complete as well.
@@
  • $T^{{\mathfrak{B}}}=\{(1,0),(1,1)\}$.
We argue that ${\mathfrak{B}}$ is a dual of $\mu$. Clearly, ${\mathfrak{B}}\not\models\mu$. Let ${\mathfrak{A}}$ be a finite $\tau$-structure with the domain $A$ such that ${\mathfrak{A}}\not\models\mu$. Let ${\mathfrak{A}}^{\prime}$ be the directed graph with the vertex set $A$ and the edge relation $E$ where $E=R^{\mathfrak{A}}\cup S^{{\mathfrak{A}}}\cup T^{{\mathfrak{A}}}$. We define $h\colon A\to\{0,1\}$ by setting $h(a)=0$ if there exists $b,c\in A$ such that $(b,c)\in R^{{\mathfrak{A}}}$ and there is a directed path from $c$ to $a$ in ${\mathfrak{A}}^{\prime}$, and $h(a)=1$ otherwise. It is straightforward to verify that $h$ is a homomorphism from ${\mathfrak{A}}$ to ${\mathfrak{B}}$.
@@ 8.4. Infinite Duals
By Properties (1) and (6) of Theorem 8.18, ${\mathfrak{B}}_{\mu}$ is always a reduct of a finitely bounded homogeneous structure. For short, we write $\Gamma_{\mu}$ for $\Gamma({\mathfrak{B}}_{\mu},\emptyset)$ and $\Gamma_{\mu,\sigma}$ for $\Gamma({\mathfrak{B}}_{\mu},\sigma)$, see Definition 8.10. For some queries $\mu$, the structure ${\mathfrak{B}}_{\mu}$ can be replaced by a simpler structure ${\mathfrak{C}}_{\mu}$. This will be convenient for some examples that we consider later, because the structure ${\mathfrak{C}}_{\mu}$ is finitely bounded and homogeneous itself. To define the respective class of queries, we need the following definition. The Gaifman graph of a relational structure ${\mathfrak{A}}$ is the undirected graph with vertex set $A$ where $a,b\in A$ are adjacent if and only if $a\neq b$ and there exists a tuple in a relation of ${\mathfrak{A}}$ that contains both $a$ and $b$. The Gaifman graph of a conjunctive query is the Gaifman graph of the canonical database of that query.
@@
For every union $\mu$ of connected conjunctive queries over a finite relational signature $\tau$ such that the Gaifman graph of each of the conjunctive queries in $\mu$ is complete, there exists a countable $\tau$-structure ${\mathfrak{C}}_{\mu}$ such that the following statements hold:
@@
    $\operatorname{Age}({\mathfrak{C}}_{\mu})$ is the class of all finite structures ${\mathfrak{A}}$ that satisfy $\neg\mu$.
Moreover, ${\mathfrak{C}}_{\mu}$ is finitely bounded, $\operatorname{Aut}({\mathfrak{C}}_{\mu})$ is oligomorphic, and a countable $\tau$-structure satisfies $\neg\mu$ if and only if it embeds into ${\mathfrak{C}}_{\mu}$.
@@
Let ${\mathfrak{A}}_{1}$ and ${\mathfrak{A}}_{2}$ be finite $\tau$-structures that satisfy $\neg\mu$ such that the substructure induced by $A_{1}\cap A_{2}$ in ${\mathfrak{A}}_{1}$ and ${\mathfrak{A}}_{2}$ is the same. Since the Gaifman graph of each of the conjunctive queries in $\mu$ is complete, the union of the structures ${\mathfrak{A}}_{1}$ and ${\mathfrak{A}}_{2}$ satisfies $\neg\mu$ as well. By Fraïssé’s Theorem (see, e.g., (Hodges, 1997)) there is a countable homogeneous $\tau$-structure ${\mathfrak{C}}_{\mu}$ such that $\operatorname{Age}({\mathfrak{C}}_{\mu})$ is the class of all finite structures that satisfy $\neg\mu$; this shows that ${\mathfrak{C}}_{\mu}$ is finitely bounded. Homogeneous structures with finite relational signature clearly have an oligomorphic automorphism group. For the final statement, see (Bodirsky, 2021, Lemma 4.1.7). ∎
Note that ${\mathfrak{C}}_{\mu}$ is homomorphically equivalent to ${\mathfrak{B}}_{\mu}$ by (Bodirsky, 2021, Lemma 4.1.7). Therefore, $\Gamma({\mathfrak{C}}_{\mu},\sigma)$ is homomorphically equivalent to $\Gamma_{\mu,\sigma}$ for any $\sigma\subseteq\tau$.
The following proposition follows straightforwardly from the definitions and provides a valued constraint satisfaction problem that is polynomial-time equivalent to the resilience problem for $\mu$, similar to Theorem 8.11.
@@
To prove a polynomial-time reduction in the other direction, let $\phi$ be a $\tau$-expression. We construct a bag database ${\mathfrak{A}}$ with signature $\tau$. The domain of ${\mathfrak{A}}$ are the variables that appear in $\phi$ and for every $S\in\tau$, we put a tuple $(x_{1},\dots,x_{n})$ in $S^{{\mathfrak{A}}}$ with a multiplicity equal to the number of occurences of $S(x_{1},\dots x_{n})$ as a summand of $\phi$. The relations $S^{{\mathfrak{A}}}$ with $S\in\sigma$ are exogenous in ${\mathfrak{A}}$, the remaining ones are endogenous. Again, ${\mathfrak{A}}$ can be computed in polynomial time and the resilience of ${\mathfrak{A}}$ with respect to $\mu$ is at most $u$ if and only if $(\phi,u)$ has a solution over $\Gamma({\mathfrak{B}},\sigma)$. ∎
In (Makhija and Gatterbauer, 2023) one may find a seemingly more general notion of exogenous tuples, where in a single relation there might be both endogenous and exogenous tuples. However, one can show that classifying the complexity of resilience problems according to our original definition also entails a classification of this variant, using the operator $\operatorname{Opt}$ and a similar reduction as in Proposition 8.20.
@@
Consider a union $\mu$ of conjunctive queries with the signature $\tau$, let $\sigma\subseteq\tau$, and let $\rho\subseteq\tau\setminus\sigma$. Suppose we would like to model the resilience problem for $\mu$ where the relations in $\sigma$ are exogenous and the relations in $\rho$ might contain both endogenous and exogenous tuples. Let ${\mathfrak{B}}$ be a dual of $\mu$ and $\Gamma$ be the expansion of $\Gamma({\mathfrak{B}},\sigma)$ where for every relational symbol $R\in\rho$, there is also a relation $(R^{x})^{\Gamma}=R^{{\mathfrak{B}}}$, i.e., a classical relation that takes values 00 and $\infty$. The resilience problem for $\mu$ with exogenous tuples specified as above is polynomial-time equivalent to $\operatorname{VCSP}(\Gamma)$ by analogous reductions as in Proposition 8.20. Note that $(R^{x})^{\Gamma}=\operatorname{Opt}\big{(}R^{\Gamma({\mathfrak{B}},\sigma)}\big{)}$ for every $R\in\rho$, and therefore by Lemma 4.6, $\operatorname{VCSP}(\Gamma)$ is polynomial-time equivalent to $\operatorname{VCSP}(\Gamma({\mathfrak{B}},\sigma))$ and thus to the resilience problem for $\mu$ where the relations in $\sigma$ are exogeneous and the relations in $\tau\setminus\sigma$ are purely endogeneous. This justifies the restriction to our setting for exogenous tuples. Moreover, the same argument shows that if resilience of $\mu$ with all tuples endogenous is in P, then all variants of resilience of $\mu$ with exogenous tuples are in P as well.
@@
We revisit Example 8.3. Consider the conjunctive query $\exists x,y,z\,(R(x,y)\wedge S(x,y,z))$ over the signature $\tau=\{R,S\}$. Note that the Gaifman graph of $\mu$ is complete; let ${\mathfrak{C}}_{\mu}$ be the structure from Theorem 8.19 and recall that it is finitely bounded and homogeneous. We construct a binary fractional polymorphism of $\Gamma({\mathfrak{C}}_{\mu},\emptyset)$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}(\Gamma({\mathfrak{C}}_{\mu},\emptyset))=\operatorname{Aut}({\mathfrak{C}}_{\mu})$. Let ${\mathfrak{M}}$ be the $\tau$-structure with domain $(C_{\mu})^{2}$ and where
  • $((b^{1}_{1},b^{2}_{1}),(b^{1}_{2},b^{2}_{2}))\in R^{{\mathfrak{M}}}$ if $(b^{1}_{1},b^{1}_{2})\in R^{{\mathfrak{C}}_{\mu}}$ and $(b^{2}_{1},b^{2}_{2})\in R^{{\mathfrak{C}}_{\mu}}$
  • $((b^{1}_{1},b^{2}_{1}),(b^{1}_{2},b^{2}_{2}),(b^{1}_{3},b^{2}_{3}))\in S^{{\mathfrak{M}}}$ if $(b^{1}_{1},b^{1}_{2},b^{1}_{3})\in S^{{\mathfrak{C}}_{\mu}}$ or $(b^{2}_{1},b^{2}_{2},b^{2}_{3})\in S^{{\mathfrak{C}}_{\mu}}$.
Similarly, let ${\mathfrak{N}}$ be the $\tau$-structure with domain $(C_{\mu})^{2}$ and where
  • $((b^{1}_{1},b^{2}_{1}),(b^{1}_{2},b^{2}_{2}))\in R^{{\mathfrak{N}}}$ if $(b^{1}_{1},b^{1}_{2})\in R^{{\mathfrak{C}}_{\mu}}$ or $(b^{2}_{1},b^{2}_{2})\in R^{{\mathfrak{C}}_{\mu}}$,
  • $((b^{1}_{1},b^{2}_{1}),(b^{1}_{2},b^{2}_{2}),(b^{1}_{3},b^{2}_{3}))\in S^{{\mathfrak{N}}}$ if $(b^{1}_{1},b^{1}_{2},b^{1}_{3})\in S^{{\mathfrak{C}}_{\mu}}$ and $(b^{2}_{1},b^{2}_{2},b^{2}_{3})\in S^{{\mathfrak{C}}_{\mu}}$.
Note that ${\mathfrak{M}}\not\models\mu$ and ${\mathfrak{N}}\not\models\mu$ and hence there are embeddings $f\colon{\mathfrak{M}}\to{\mathfrak{C}}_{\mu}$ and $g\colon{\mathfrak{N}}\to{\mathfrak{C}}_{\mu}$. Clearly, both $f$ and $g$ regarded as operations on the set $C_{\mu}$ are pseudo cyclic (but in general not cyclic) and canonical with respect to $\operatorname{Aut}({\mathfrak{C}}_{\mu})$ (see Claim 6 in Proposition 8.27 for a detailed argument of this type). Let $\omega$ be the fractional operation given by $\omega(f)=\frac{1}{2}$ and $\omega(g)=\frac{1}{2}$. Then $\omega$ is a binary fractional polymorphism of $\Gamma:=\Gamma({\mathfrak{C}}_{\mu},\emptyset)$: for $b^{1},b^{2}\in(C_{\mu})^{2}$ we have
@@
Let ${\mathfrak{C}}_{\mu}$ be the homogeneous dual of $\mu$ that embeds every countable structure ${\mathfrak{A}}$ that does not satisfy $\mu$. Let $C$ be the domain of ${\mathfrak{C}}_{\mu}$ and let $\Gamma:=\Gamma({\mathfrak{C}}_{\mu},\emptyset)$. We show that $\Gamma$ has a ternary canonical pseudo cyclic fractional polymorphism (with respect to $\operatorname{Aut}({\mathfrak{C}}_{\mu})$), which implies tractability of $\operatorname{VCSP}(\Gamma)$ (Theorem 7.18) and hence of the resilience problem for $\mu$ (Proposition 8.20). To increase readability, we write $R$ for $R^{{\mathfrak{C}}_{\mu}}$ and write $\breve{R}$ for $\{(a,b)\mid(b,a)\in R^{{\mathfrak{C}}_{\mu}}\}$. Note that since ${\mathfrak{C}}_{\mu}\models\neg\mu$, for every $a,b\in C$, we have precisely one of the following: $(a,b)\in R$, $(a,b)\in\breve{R}$, or $(a,b)\notin R\cup\breve{R}$. Let ${\mathfrak{M}}$ be an $\{R\}$-structure with the domain $C^{3}$ such that
@@
Note that items (2)-(4) are just cyclic shifts of the same condition. It is straightforward to verify that ${\mathfrak{M}}\models\neg\mu$. For example, if ${\mathfrak{M}}\models R((x,y,z),(u,v,w))$ because of item (1), then at least two of $(u,x),(v,y),(w,z)$ lie in $\breve{R}$ and hence ${\mathfrak{M}}\models\neg R((u,v,w),(x,y,z))$ by definition. Therefore, there is an embedding $f$ of ${\mathfrak{M}}$ into ${\mathfrak{C}}_{\mu}$. By the definition of ${\mathfrak{M}}$, the operation $f$ is pseudo cyclic and canonical with respect to $\operatorname{Aut}({\mathfrak{C}}_{\mu})$ (this is easy to see, because ${\mathfrak{C}}_{\mu}$ is homogeneous; a detailed argument of this type can be found in Claim 6 in Proposition 8.27). The idea is that $f$ has the behavior of a majority operation on orbits of pairs of distinct elements.

### v6

@@ 8.3. Regular Path Queries
We introduce two-way regular path queries and show that they have finite duals. A two-way regular path query (2RPQ) over a finite relational signature $\tau$ is a regular expression over the alphabet $\Sigma_{\tau}$ that consists of the symbols $R$ and $R^{-}$ for every binary relation symbol $R\in\tau$. We assume that regular expressions are built up from alphabet symbols, $\emptyset$, $\varepsilon$, ++ (union), ; (concatenation), and $\cdot^{*}$ (Kleene star). As usual, we may omit the concatenation symbol for brevity.
@@
We consider a Boolean version of the resilience problem for 2RPQs. For a 2RPQ $\mu$ over a signature $\tau$ and a set $\sigma\subseteq\tau$ of exogenous relations, this problem is as follows: given a finite (bag) database $\mathfrak{A}$} and a $k\geq 1$, decide whether it suffices to remove at most $k$ tuples from $\mathfrak{A}$} to achieve that $\mu({\mathfrak{A}})=\emptyset$. It is folklore that given a 2RPQ $\mu$ and a database $\mathfrak{A}$}, one can decide in polynomial time whether $\mu({\mathfrak{A}})=\emptyset$, e.g. by forming the product of $\mathfrak{A}$} and a finite two-way automaton $A$ with $L(A)=L(\mu)$ and then checking whether some element with a final state in the second component is reachable from any element with the initial state in the second component. Consequently, for any 2RPQ $\mu$ with exogenous relations from any set $\sigma$ the resilience problem is in NP by a straightforward guess-and-check algorithm.
@@
Let $\mu$ be a 2RPQ over a relational signature $\tau$ and let $\sigma\subseteq\tau$. Then the resilience problem for $\mu$ with exogenous relations from $\sigma$ is in P or NP-complete. Moreover, it is decidable whether the resilience problem for a given 2RPQ $\mu$ and set of exogenous relations $\sigma$ is in P.
@@
Consider the 2RPQ $\mu$ over the signature $\{R\}$ given by $R(RR^{-})^{*}RR$. It is well-known (see Propositions 1.6 and 1.13 in (Hell and Nešetřil, 2004)) that the relational $\{R\}$-structure $\mathfrak{B}$} with domain $\{0,1,2\}$ and the relation $R^{{\mathfrak{B}}}:=\{(0,1),(1,2)\}$ is a dual of $\mu$. Let $\Gamma:=\Gamma({\mathfrak{B}};\emptyset)$. Note that the crisp unary relation $\{0,1\}$ is pp-expressible in $\Gamma$ by
@@
Let $R^{\prime}(x,y):=R(x,y)+\phi(x)+\phi(y)$ and note that $R^{\prime}\in\langle\Gamma\rangle$. It is easy to see that the valued structure $(\{0,1,2\};R^{\prime})$ is fractionally homomorphically equivalent to the valued structure $\Gamma_{<}$ from Example 2.1. In other words, $\Gamma$ $p$-constructs $\Gamma_{<}$. Since $\operatorname{VCSP}(\Gamma_{<})$ is NP-complete (Example 4.6), $\operatorname{VCSP}(\Gamma)$ is NP-complete by Lemma 5.16 (every finite-domain VCSP is in NP). Therefore, the resilience problem for the 2RPQ $\mu$ is NP-complete as well.
@@
  • $T^{{\mathfrak{B}}}=\{(1,0),(1,1)\}$
(see Figure 3). We argue that $\mathfrak{B}$} is a dual of $\mu$. Clearly, ${\mathfrak{B}}\not\models\mu$. Let $\mathfrak{A}$} be a finite $\tau$-structure with the domain $A$ such that ${\mathfrak{A}}\not\models\mu$. Let $\mathfrak{A}$$^{\prime}$ be the directed graph with the vertex set $A$ and the edge relation $E$ where $E=R^{\mathfrak{A}}\cup S^{{\mathfrak{A}}}\cup T^{{\mathfrak{A}}}$. We define $h\colon A\to\{0,1\}$ by setting $h(a)=0$ if there exists $b,c\in A$ such that $(b,c)\in R^{{\mathfrak{A}}}$ and there is a directed path from $c$ to $a$ in $\mathfrak{A}$$^{\prime}$, and $h(a)=1$ otherwise. It is straightforward to verify that $h$ is a homomorphism from $\mathfrak{A}$} to $\mathfrak{B}$}.
@@
Figure 3. The dual structure $\mathfrak{B}$} from Example 8.17.:
@@ 8.4. Infinite Duals
By Properties (1) and (6) of Theorem 8.18, ${\mathfrak{B}}_{\mu}$ is always a reduct of a finitely bounded homogeneous structure. For short, we write $\Gamma_{\mu}$ for $\Gamma({\mathfrak{B}}_{\mu},\emptyset)$ and $\Gamma_{\mu,\sigma}$ for $\Gamma({\mathfrak{B}}_{\mu},\sigma)$, see Definition 8.10. For some queries $\mu$, the structure ${\mathfrak{B}}_{\mu}$ can be replaced by a simpler structure ${\mathfrak{H}}_{\mu}$. This will be convenient for some examples that we consider later, because the structure ${\mathfrak{H}}_{\mu}$ is finitely bounded and homogeneous itself. To define the respective class of queries, we need the following definition. The Gaifman graph of a relational structure $\mathfrak{A}$} is the undirected graph with vertex set $A$ where $a,b\in A$ are adjacent if and only if $a\neq b$ and there exists a tuple in a relation of $\mathfrak{A}$} that contains both $a$ and $b$. The Gaifman graph of a conjunctive query is the Gaifman graph of the canonical database of that query.
@@
For every union $\mu$ of conjunctive queries over a finite relational signature $\tau$ such that the Gaifman graph of each of the conjunctive queries in $\mu$ is complete, there exists a countable $\tau$-structure ${\mathfrak{H}}_{\mu}$ such that the following statements hold:
@@
    $\operatorname{Age}({\mathfrak{H}}_{\mu})$ is the class of all finite structures $\mathfrak{A}$} that satisfy $\neg\mu$.
Moreover, ${\mathfrak{H}}_{\mu}$ is finitely bounded, $\operatorname{Aut}({\mathfrak{H}}_{\mu})$ is oligomorphic, and a countable $\tau$-structure satisfies $\neg\mu$ if and only if it embeds into ${\mathfrak{H}}_{\mu}$.
@@
Let ${\mathfrak{A}}_{1}$ and ${\mathfrak{A}}_{2}$ be finite $\tau$-structures that satisfy $\neg\mu$ such that the substructure induced by $A_{1}\cap A_{2}$ in ${\mathfrak{A}}_{1}$ and ${\mathfrak{A}}_{2}$ is the same. Since the Gaifman graph of each of the conjunctive queries in $\mu$ is complete, the union of the structures ${\mathfrak{A}}_{1}$ and ${\mathfrak{A}}_{2}$ satisfies $\neg\mu$ as well, because there are no relations between elements from $A_{1}\setminus A_{2}$ and $A_{2}\setminus A_{1}$ in the union. By Fraïssé’s Theorem (see, e.g., (Hodges, 1997)) there is a countable homogeneous $\tau$-structure ${\mathfrak{H}}_{\mu}$ such that $\operatorname{Age}({\mathfrak{H}}_{\mu})$ is the class of all finite structures that satisfy $\neg\mu$; this shows that ${\mathfrak{H}}_{\mu}$ is finitely bounded. Homogeneous structures with finite relational signature clearly have an oligomorphic automorphism group. For the final statement, see (Bodirsky, 2021, Lemma 4.1.7). ∎
Note that ${\mathfrak{H}}_{\mu}$ is homomorphically equivalent to ${\mathfrak{B}}_{\mu}$ by (Bodirsky, 2021, Lemma 4.1.7). Therefore, $\Gamma({\mathfrak{H}}_{\mu},\sigma)$ is homomorphically equivalent to $\Gamma_{\mu,\sigma}$ for any $\sigma\subseteq\tau$.
We now continue with general propositions that apply to all unions of conjunctive queries $\mu$. The following proposition follows straightforwardly from the definitions and provides a valued constraint satisfaction problem that is polynomial-time equivalent to the resilience problem for $\mu$, similar to Theorem 8.11.
@@
To prove a polynomial-time reduction in the other direction, let $\phi$ be a $\tau$-expression. We construct a bag database $\mathfrak{A}$} with signature $\tau$. The domain of $\mathfrak{A}$} are the variables that appear in $\phi$ and for every $S\in\tau$, we put a tuple $(x_{1},\dots,x_{n})$ in $S^{{\mathfrak{A}}}$ with a multiplicity equal to the number of occurrences of $S(x_{1},\dots x_{n})$ as a summand of $\phi$. The relations $S^{{\mathfrak{A}}}$ with $S\in\sigma$ are exogenous in $\mathfrak{A}$}, the remaining ones are endogenous. Again, $\mathfrak{A}$} can be computed in polynomial time and the resilience of $\mathfrak{A}$} with respect to $\mu$ is at most $u$ if and only if $(\phi,u)$ has a solution over $\Gamma({\mathfrak{B}},\sigma)$. ∎
@@
The reason why we restrict to study of resilience problems in bag semantics and do not consider resilience problems in set semantics comes from the proof of Proposition 8.20. To model resilience problems in set semantics, we would have to restrict the instances of the corresponding VCSP to instances without repeating summands. Little is known about the complexity of this variant of VCSPs and there is no algebraic theory describing the complexity of such problems (Osipov, 2025). Consequently, the theory of constraint satisfaction problems does not provide tools to classify the complexity of resilience problems in set semantics; nevertheless, tractability results in bag semantics imply tractability results for the same queries in set semantics.
In (Makhija and Gatterbauer, 2023) one may find a seemingly more general notion of exogenous tuples, where in a single relation there might be both endogenous and exogenous tuples. However, one can show that classifying the complexity of resilience problems according to our original definition also entails a classification of this variant, using the operator $\operatorname{Opt}$ and a similar reduction as in Proposition 8.20, as explained in the following remark.
Remark 8.22.
Consider a union $\mu$ of conjunctive queries with the signature $\tau$, let $\sigma\subseteq\tau$, and let $\rho\subseteq\tau\setminus\sigma$. Suppose we would like to model the resilience problem for $\mu$ where the relations in $\sigma$ are exogenous and the relations in $\rho$ might contain both endogenous and exogenous tuples. Let $\mathfrak{B}$} be a dual of $\mu$ and $\Gamma$ be the expansion of $\Gamma({\mathfrak{B}},\sigma)$ where for every relational symbol $R\in\rho$, there is also a relation $(R^{x})^{\Gamma}=R^{{\mathfrak{B}}}$, i.e., a classical relation that takes values 0 and $\infty$. The resilience problem for $\mu$ with exogenous tuples specified as above is polynomial-time equivalent to $\operatorname{VCSP}(\Gamma)$ by analogous reductions as in Proposition 8.20. Note that $(R^{x})^{\Gamma}=\operatorname{Opt}\big(R^{\Gamma({\mathfrak{B}},\sigma)}\big)$ for every $R\in\rho$, and therefore by Lemma 4.5, $\operatorname{VCSP}(\Gamma)$ is polynomial-time equivalent to $\operatorname{VCSP}(\Gamma({\mathfrak{B}},\sigma))$ and thus to the resilience problem for $\mu$ where the relations in $\sigma$ are exogenous and the relations in $\tau\setminus\sigma$ are purely endogenous. This justifies the restriction to our setting for exogenous tuples. Moreover, the same argument shows that if resilience of $\mu$ with all tuples endogenous is in P, then all variants of resilience of $\mu$ with exogenous tuples are in P as well.
@@
Remark 8.23.
Note that if the multiplicities in the given database are stored in binary, then this corresponds to binary representations of multiplicities of summands in instances of the VCSP, and as we have mentioned earlier (Remark 7.16 and Remark 7.20) many of the algorithmic results about VCSPs apply even in this setting.
@@
We revisit Example 8.3. Consider the conjunctive query $\exists x,y,z\,(R(x,y)\wedge S(x,y,z))$ over the signature $\tau=\{R,S\}$. We observe that its resilience problem is in P. If $\mathfrak{A}$} is the input database, let $m$ be the multiplicity of $(a,b)$ in $R^{\mathfrak{A}}$ and let $m^{\prime}$ be the sum of multiplicities of tuples of the form $(a,b,c)$ for some $c\in A$ in $S^{\mathfrak{A}}$. Then, if $m\leq m^{\prime}$, we remove all copies of $(a,b)$ from $R^{\mathfrak{A}}$ and otherwise we remove all tuples of the form $(a,b,c)$ from $S^{\mathfrak{A}}$.
Note that the Gaifman graph of $\mu$ is complete; let ${\mathfrak{H}}_{\mu}$ be the structure from Theorem 8.19 and recall that it is finitely bounded and homogeneous. We construct a binary fractional polymorphism of $\Gamma({\mathfrak{H}}_{\mu},\emptyset)$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}(\Gamma({\mathfrak{H}}_{\mu},\emptyset))=\operatorname{Aut}({\mathfrak{H}}_{\mu})$. Let $\mathfrak{M}$} be the $\tau$-structure with domain $(H_{\mu})^{2}$ and where
  • $((b^{1}_{1},b^{2}_{1}),(b^{1}_{2},b^{2}_{2}))\in R^{{\mathfrak{M}}}$ if $(b^{1}_{1},b^{1}_{2})\in R^{{\mathfrak{H}}_{\mu}}$ and $(b^{2}_{1},b^{2}_{2})\in R^{{\mathfrak{H}}_{\mu}}$
  • $((b^{1}_{1},b^{2}_{1}),(b^{1}_{2},b^{2}_{2}),(b^{1}_{3},b^{2}_{3}))\in S^{{\mathfrak{M}}}$ if $(b^{1}_{1},b^{1}_{2},b^{1}_{3})\in S^{{\mathfrak{H}}_{\mu}}$ or $(b^{2}_{1},b^{2}_{2},b^{2}_{3})\in S^{{\mathfrak{H}}_{\mu}}$.
Similarly, let $\mathfrak{N}$} be the $\tau$-structure with domain $(H_{\mu})^{2}$ and where
  • $((b^{1}_{1},b^{2}_{1}),(b^{1}_{2},b^{2}_{2}))\in R^{{\mathfrak{N}}}$ if $(b^{1}_{1},b^{1}_{2})\in R^{{\mathfrak{H}}_{\mu}}$ or $(b^{2}_{1},b^{2}_{2})\in R^{{\mathfrak{H}}_{\mu}}$,
  • $((b^{1}_{1},b^{2}_{1}),(b^{1}_{2},b^{2}_{2}),(b^{1}_{3},b^{2}_{3}))\in S^{{\mathfrak{N}}}$ if $(b^{1}_{1},b^{1}_{2},b^{1}_{3})\in S^{{\mathfrak{H}}_{\mu}}$ and $(b^{2}_{1},b^{2}_{2},b^{2}_{3})\in S^{{\mathfrak{H}}_{\mu}}$.
Note that ${\mathfrak{M}}\not\models\mu$ and ${\mathfrak{N}}\not\models\mu$ and hence there are embeddings $f\colon{\mathfrak{M}}\to{\mathfrak{H}}_{\mu}$ and $g\colon{\mathfrak{N}}\to{\mathfrak{H}}_{\mu}$. Clearly, both $f$ and $g$ regarded as operations on the set $H_{\mu}$ are pseudo cyclic (but in general not cyclic) and canonical with respect to $\operatorname{Aut}({\mathfrak{H}}_{\mu})$ (see Claim 6 in Proposition 8.29 for a detailed argument of this type). Let $\omega$ be the fractional operation given by $\omega(f)=\frac{1}{2}$ and $\omega(g)=\frac{1}{2}$. Then $\omega$ is a binary fractional polymorphism of $\Gamma:=\Gamma({\mathfrak{H}}_{\mu},\emptyset)$: for $b^{1},b^{2}\in(H_{\mu})^{2}$ we have
@@
Let ${\mathfrak{H}}_{\mu}$ be the homogeneous dual of $\mu$ that embeds every countable structure $\mathfrak{A}$} that does not satisfy $\mu$. Let let $\Gamma:=\Gamma({\mathfrak{H}}_{\mu},\emptyset)$ and let $C$ be the domain of $\Gamma$. We show that $\Gamma$ has a ternary canonical pseudo cyclic fractional polymorphism (with respect to $\operatorname{Aut}({\mathfrak{H}}_{\mu})$), which implies tractability of $\operatorname{VCSP}(\Gamma)$ (Theorem 7.19) and hence of the resilience problem for $\mu$ (Proposition 8.20). To increase readability, we write $R$ for $R^{{\mathfrak{H}}_{\mu}}$ and write $\breve{R}$ for $\{(a,b)\mid(b,a)\in R^{{\mathfrak{H}}_{\mu}}\}$. Note that since ${\mathfrak{H}}_{\mu}\models\neg\mu$, for every $a,b\in C$, we have precisely one of the following: $(a,b)\in R$, $(a,b)\in\breve{R}$, or $(a,b)\notin R\cup\breve{R}$. Let $\mathfrak{M}$} be an $\{R\}$-structure with the domain $C^{3}$ such that
@@
Note that items (2)-(4) are just cyclic shifts of the same condition. It is straightforward to verify that ${\mathfrak{M}}\models\neg\mu$. For example, if ${\mathfrak{M}}\models R((x,y,z),(u,v,w))$ because of item (1), then at least two of $(u,x),(v,y),(w,z)$ lie in $\breve{R}$ and hence ${\mathfrak{M}}\models\neg R((u,v,w),(x,y,z))$ by definition. Therefore, there is an embedding $f$ of $\mathfrak{M}$} into ${\mathfrak{H}}_{\mu}$. By the definition of $\mathfrak{M}$}, the operation $f$ is pseudo cyclic and canonical with respect to $\operatorname{Aut}({\mathfrak{H}}_{\mu})$ (this is easy to see, because ${\mathfrak{H}}_{\mu}$ is homogeneous; a detailed argument of this type can be found in Claim 6 in Proposition 8.29). The idea is that $f$ has the behavior of a majority operation on orbits of pairs of distinct elements.